OpenStudy (anonymous):

Write the equation of the line that passes through (–2, 1) and is perpendicular to the line 3x – 2y = 5.

OpenStudy (texaschic101):

first we need to find the slope in the equation 3x - 2y = 5. Do you know how to do that ?

OpenStudy (texaschic101):

put the equation in y = mx + b form, and m will be the slope

OpenStudy (anonymous):

it would be y= 3/2x-5/2 right?

OpenStudy (texaschic101):

3x - 2y = 5 -2y = -3x + 5 y = 3/2x - 5/2 yes...the slope is 3/2, but we need a perpendicular line, therefore, we need the negative reciprocal of that slope. All that means is " flip " the slope and change the sign. Therefore, the slope we need to use is -2/3. Do you see how I flipped the slope and changed the sign ? Do you understand so far ?

OpenStudy (anonymous):

yepp

OpenStudy (texaschic101):

now we can use either y = mx + b form or y - y1 = m(x - x1) form...you will arrive at the same answer either one you choose to use. Which formula do you want to use ?

OpenStudy (texaschic101):

or I can show you how to use both ....do you want me to show you both ?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope(m) = -2/3 (-2,1) x1 = -2 and y1 = 1 now we sub y - 1 = -2/3(x - (-2) y - 1 = -2/3(x + 2) y - 1 = -2/3x - 4/3 y = -2/3x - 4/3 + 1 y = -2/3x - 4/3 + 3/3 y = -2/3x - 1/3 <-- your perpendicular line any questions ?

OpenStudy (anonymous):

No. Thank you !

OpenStudy (texaschic101):

you could have also done this.. y = mx + b and find b slope(m) = -2/3 (-2,1) x = -2 and y = 1 now we sub y = mx + b 1 = -2/3(-2) + b 1 = 4/3 + b 1 - 4/3 = b 3/3 - 4/3 = b -1/3 = b leaving you with y = -2/3x - 1/3

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

anytime :)